Let f(z)=1/z^5 . Use the polar form of the Cauchy Riemann equations to determine where f is differentiable

1
Expert's answer
2022-05-08T14:45:48-0400

"z=re^{i\theta}," where "r>0" and "\theta\in[0,2\pi)". This is the polar form of the complex variable. "f=r^{-5}e^{-5i\theta}." Remind that "e^{i\theta}=cos(\theta)+i\,sin(\theta)". We receive: "f=r^{-5}(cos(5\theta)-i\,sin(5\theta))=u+iv," where "u=r^{-5}cos(5\theta)", "v=-r^{-5}sin(5\theta)". Cauchy-Riemann equations in polar coordinates have the following form: "\frac{\partial u}{\partial r}=\frac{1}{r}\frac{\partial v}{\partial\theta}", "\frac{\partial v}{\partial r}=-\frac1r\frac{\partial u}{\partial\theta}". "\frac{\partial u}{\partial r}=-5r^{-6}cos(5\theta)," "\frac{\partial u}{\partial \theta}=-5r^{-5}sin(5\theta)", "\frac{\partial v}{\partial r}=5r^{-6}sin(5\theta)," "\frac{\partial v}{\partial\theta}=-5r^{-5}cos(5\theta)". As we can see, the Cauchy-Riemann equations are satisfied at all points, except "r=0" . Functions "u,v" are continuously differentiable at all points, except "r=0". Thus, due to the respective Theorem from complex analysis, the function "f" is differentiable at all points except .